### v1

@@ C.4 Non-Parametric Causal Analysis
C.4.1 Distribution of Estimated Individual Treatment Effects
@@
C.4.2 Treatment Effects by Working Years
@@
C.4.3 Treatment Effects by Research Productivity
@@ D.1 Non-Causal Analysis
(Intercept) 4.7709 0.0060 792.959 $<$2e-16***
Gender (Female) -0.0400 0.0027 -14.875 $<$2e-16***
@@
Titles: Associate Professor 0.0654 0.0034 19.358 $<$2e-16***
Titles: Professor 0.1812 0.0040 44.783 $<$2e-16***
University Code (DRU(H)) 0.0534 0.0040 13.409 $<$2e-16***
University Code (DU/VA) 0.1653 0.0041 39.944 $<$2e-16***
Department Code (B) 0.2603 0.0058 44.737 $<$2e-16***
Department Code (MHS) 0.2836 0.0041 69.441 $<$2e-16***
Department Code (NS) 0.0715 0.0045 15.923 $<$2e-16***
Department Code (SS) 0.0720 0.0043 16.663 $<$2e-16***
Department Code (TE) 0.0892 0.0046 19.461 $<$2e-16***
Working Years -0.0021 0.0002 -12.545 $<$2e-16***
@@ D.2 Parametric Causal Analysis
Distance Distance 0.8043 -0.0001
Google Scholar ID Yes Binary -0.1162 -0.2857
Titles: Assistant Professor Binary 0.1258 0.0181
Titles: Associate Professor Binary 0.0178 -0.0347
Titles: Professor Binary -0.1436 0.0166
Working Years Continuous -0.2927 -0.0833
University Code: Bachelor/Master Binary 0.0052 0.0029
University Code: DRU(H) Binary -0.0037 -0.0100
University Code: DU/VA Binary -0.0015 -0.0130
Department Code: AH Binary 0.0119 0.0288
Department Code: B Binary -0.0273 -0.0261
Department Code: MHS Binary 0.0967 0.0487
Department Code: NS Binary -0.0541 -0.0217
Department Code: SS Binary 0.0543 -0.0108
Department Code: TE Binary -0.0814 -0.0190
$\log_{10}$(i10index) Continuous -0.6261 0.2676
@@
(Intercept) 5.0194 0.0075 665.570 $<$2e-16***
Gender (Female) -0.0293 0.0040 -7.346 2.26e-13***
Google Scholar ID (Yes) 0.0029 0.0036 0.798 0.4251
Assistant Prof $\times$ Working Years -0.0071 0.0004 -18.011 $<$2e-16***
Associate Prof $\times$ Working Years -0.0017 0.0003 -6.580 5.01e-11***
Professor $\times$ Working Years 0.0024 0.0002 12.311 $<$2e-16***
bachelor/master $\times$ AH $\times$ log10(i10-index) -0.0971 0.0092 -10.524 $<$2e-16***
DRU(H) $\times$ AH $\times$ log10(i10-index) -0.0900 0.0068 -13.257 $<$2e-16***
DU/VA $\times$ AH $\times$ log10(i10-index) -0.0517 0.0075 -6.864 7.27e-12***
bachelor/master $\times$ B $\times$ log10(i10-index) 0.0260 0.0144 1.805 0.0712
DRU(H) $\times$ B $\times$ log10(i10-index) 0.1082 0.0087 12.367 $<$2e-16***
DU/VA $\times$ B $\times$ log10(i10-index) 0.1791 0.0101 17.786 $<$2e-16***
bachelor/master $\times$ MHS $\times$ log10(i10-index) -0.0253 0.0117 -2.159 0.0309*
DRU(H) $\times$ MHS $\times$ log10(i10-index) 0.0971 0.0063 15.536 $<$2e-16***
DU/VA $\times$ MHS $\times$ log10(i10-index) 0.1783 0.0046 38.502 $<$2e-16***
bachelor/master $\times$ NS $\times$ log10(i10-index) -0.0757 0.0115 -6.573 5.29e-11***
DRU(H) $\times$ NS $\times$ log10(i10-index) -0.0482 0.0071 -6.810 1.06e-11***
DU/VA $\times$ NS $\times$ log10(i10-index) 0.0452 0.0053 8.582 $<$2e-16***
bachelor/master $\times$ SS $\times$ log10(i10-index) -0.0750 0.0095 -7.881 3.72e-15***
DRU(H) $\times$ SS $\times$ log10(i10-index) -0.0613 0.0068 -9.072 $<$2e-16***
DU/VA $\times$ SS $\times$ log10(i10-index) 0.0499 0.0061 8.154 4.12e-16***
bachelor/master $\times$ TE $\times$ log10(i10-index) -0.0623 0.0132 -4.731 2.28e-06***
DRU(H) $\times$ TE $\times$ log10(i10-index) -0.0084 0.0075 -1.119 0.2632
DU/VA $\times$ TE $\times$ log10(i10-index) 0.0486 0.0053 9.112 $<$2e-16***
Table 10: Three-Way Interaction Regression: Gender, Rank × Experience, and Productivity × Context Effects
@@ D.3 Non-Parametric Causal Analysis
In the full dataset, the non-parametric Causal Forest model yields an Average Treatment Effect (ATE) of $-0.0266$ ($SE=0.0038$), consistent with previous estimates. The Individual Treatment Effects (ITEs), shown in Figure D.3, range from $-0.16$ to $0.08$, confirming considerable heterogeneity in gender-based salary disparities. Subgroup analyses (Figure 14) indicate a trend toward improved gender pay equity with increasing years of experience.

### v2

@@ C.4 Non-Parametric Causal Analysis
C.4.1 Robustness Check for Different Hyperparameters
To assess robustness to hyperparameter choices, we re-estimated the causal forest model under alternative specifications of minimum node size, subsample fraction, number of covariates (mtrymtry), and number of trees. As shown in Table 8, the estimated Average Treatment Effect (ATE) remained stable across specifications, with values consistently around 0.026-0.026 and standard errors near 0.004. This indicates that our findings are not sensitive to tuning parameters and are robust to reasonable variations in model configuration.
Specification ATE Estimate Standard Error
Baseline (tuned mtrymtry, sample.fraction = 0.49) -0.0264 0.0039
Min node size = 10 -0.0264 0.0039
Min node size = 20 -0.0264 0.0039
Sample fraction = 0.45 -0.0263 0.0039
$mtry=\lceil p/3\rceil=8$ -0.0270 0.0040
$mtry=\lceil p/2\rceil=11$ -0.0266 0.0039
Number of trees = 1500 -0.0265 0.0039
Number of trees = 5000 -0.0264 0.0039
Table 8: Robustness of Causal Forest ATE Estimates to Alternative Hyperparameter Specifications
C.4.2 Distribution of Estimated Individual Treatment Effects
@@
C.4.3 Treatment Effects by Working Years
@@
C.4.4 Treatment Effects by Research Productivity
@@ D.1 Non-Causal Analysis
(Intercept) 4.7709 0.0060 792.959 <<2e-16***
Gender (Female) -0.0400 0.0027 -14.875 <<2e-16***
@@
Titles: Associate Professor 0.0654 0.0034 19.358 <<2e-16***
Titles: Professor 0.1812 0.0040 44.783 <<2e-16***
University Code (DRU(H)) 0.0534 0.0040 13.409 <<2e-16***
University Code (DU/VA) 0.1653 0.0041 39.944 <<2e-16***
Department Code (B) 0.2603 0.0058 44.737 <<2e-16***
Department Code (MHS) 0.2836 0.0041 69.441 <<2e-16***
Department Code (NS) 0.0715 0.0045 15.923 <<2e-16***
Department Code (SS) 0.0720 0.0043 16.663 <<2e-16***
Department Code (TE) 0.0892 0.0046 19.461 <<2e-16***
Working Years -0.0021 0.0002 -12.545 <<2e-16***
@@ D.2 Parametric Causal Analysis
Distance Distance 0.786 -0.000
Titles: Assistant Professor Binary 0.126 -0.000
Titles: Associate Professor Binary 0.018 -0.010
Titles: Professor Binary -0.144 0.010
Working Years Continuous -0.293 -0.039
University Code: Bachelor/Master Binary 0.005 0.015
University Code: DRU(H) Binary -0.004 -0.010
University Code: DU/VA Binary -0.002 -0.005
Department Code: AH Binary 0.012 0.037
Department Code: B Binary -0.027 -0.015
Department Code: MHS Binary 0.097 0.057
Department Code: NS Binary -0.054 -0.021
Department Code: SS Binary 0.054 -0.035
Department Code: TE Binary -0.081 -0.022
$\log_{10}$(i10index) Continuous -0.626 0.024
@@
(Intercept) 5.0186 0.0076 664.39 <<2e-16***
Gender (Female) -0.0269 0.0040 -6.74 1.7e-11***
Google Scholar ID (Yes) 0.0022 0.0037 0.61 0.5413
Assistant Prof $\times$ Working Years -0.0071 0.0004 -17.98 <<2e-16***
Associate Prof $\times$ Working Years -0.0017 0.0003 -6.45 1.2e-10***
Professor $\times$ Working Years 0.0026 0.0002 13.10 <<2e-16***
bachelor/master $\times$ AH $\times$ log10(i10-index) -0.0995 0.0093 -10.65 <<2e-16***
DRU(H) $\times$ AH $\times$ log10(i10-index) -0.0921 0.0067 -13.69 <<2e-16***
DU/VA $\times$ AH $\times$ log10(i10-index) -0.0552 0.0076 -7.27 4.1e-13***
bachelor/master $\times$ B $\times$ log10(i10-index) 0.0324 0.0149 2.18 0.0296*
DRU(H) $\times$ B $\times$ log10(i10-index) 0.1081 0.0088 12.25 <<2e-16***
DU/VA $\times$ B $\times$ log10(i10-index) 0.1760 0.0102 17.22 <<2e-16***
bachelor/master $\times$ MHS $\times$ log10(i10-index) -0.0308 0.0119 -2.58 0.0099**
DRU(H) $\times$ MHS $\times$ log10(i10-index) 0.1030 0.0062 16.60 <<2e-16***
DU/VA $\times$ MHS $\times$ log10(i10-index) 0.1758 0.0046 38.01 <<2e-16***
bachelor/master $\times$ NS $\times$ log10(i10-index) -0.0739 0.0109 -6.77 1.4e-11***
DRU(H) $\times$ NS $\times$ log10(i10-index) -0.0491 0.0072 -6.87 7.0e-12***
DU/VA $\times$ NS $\times$ log10(i10-index) 0.0445 0.0053 8.44 <<2e-16***
bachelor/master $\times$ $S$ $\times$ log10(i10-index) -0.0776 0.0095 -8.15 4.4e-16***
DRU(H) $\times$ $S$ $\times$ log10(i10-index) -0.0632 0.0068 -9.33 <<2e-16***
DU/VA $\times$ $S$ $\times$ log10(i10-index) 0.0464 0.0060 7.66 2.0e-14***
bachelor/master $\times$ TE $\times$ log10(i10-index) -0.0757 0.0146 -5.22 1.8e-07***
DRU(H) $\times$ TE $\times$ log10(i10-index) -0.0126 0.0077 -1.65 0.0986.
DU/VA $\times$ TE $\times$ log10(i10-index) 0.0453 0.0054 8.43 <<2e-16***
Table 11: Three-Way Interaction Regression: Gender, Rank $\times$ Experience, and Productivity $\times$ Context Effects
@@ D.3 Non-Parametric Causal Analysis
In the full dataset, the non-parametric Causal Forest model yields an Average Treatment Effect (ATE) of 0.0262-0.0262 ($SE=0.0036$), consistent with previous estimates. The Individual Treatment Effects (ITEs), shown in Figure 14, range from 0.066-0.066 to 0.016-0.016, confirming considerable heterogeneity in gender-based salary disparities. Subgroup analyses (Figure 15) indicate a trend toward improved gender pay equity with increasing years of experience.